# Cross-lingual Offensive Language Detection: A Systematic Review of Dataset, Approach and Challenge

### v2

@@ 7.4.6. Application of Large Language Models
With the emergence of Large Language Models (LLMs), the landscape of offensive language detection has rapidly shifted. Models such as GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), Fine-tuned LAnguage Net (FLAN) (Wei et al., 2021) have demonstrated impressive multilingual capabilities, and recent surveys have highlighted their growing importance (Zhao et al., 2023). Instruction-tuned LLMs with prompting can already perform competitive detection across monolingual and multilingual settings without the need for task-specific fine-tuning (Plaza-del arco et al., 2023). Prompt design strategies, such as incorporating task descriptions or input-label demonstrations, further enhance performance (Chen et al., 2022; Han and Tang, 2022; Roy et al., 2023). LLMs also show promise in addressing data scarcity by generating synthetic examples for low-resource languages (Hartvigsen et al., 2022), and in providing explanations that improve interpretability (Wang et al., 2023). Despite their advantages, LLMs are not a full replacement for prior methods. They often underperform specialised multilingual PLMs on fine-grained tasks when little prompting expertise is available, and they remain costly to deploy at scale. Traditional approaches, such as leveraging domain-specific lexicons (e.g., HurtLex) or designing language-family-specific PLMs (e.g., MuRIL, AraBERT), still work well for targeted low-resource or domain-specific scenarios, offering more efficiency and control. A key direction is to explore how LLMs and prior CLTL techniques can be combined. For example, lexicon features and culturally-aware embeddings can be integrated as prompts or external knowledge sources to guide LLMs, while smaller multilingual PLMs can be lightweight alternatives in resource-constrained environments.
Overall, while early work on LLMs in this area was scarce at the time of our survey design, the rapid pace of development since then underscores their central role going forward. Future studies should explicitly compare LLM-based approaches with prior CLTL techniques to establish where LLMs truly offer an advantage and where traditional methods remain more effective.
@@ 8. Conclusion
  • Cultural sensitivity: Offensive expressions are often context-dependent, and models trained on data from one cultural context may misclassify content in another. Culturally-aware annotation and evaluation with native speakers from target communities are necessary.
@@ Acknowledgements.
Aiqi Jiang is funded by China Scholarship Council (CSC Funding, No. 201908510140). Arkaitz Zubiaga acknowledges support from the European Union and UK Research and Innovation under Grant No. 101073351 as part of Marie Skłodowska-Curie Actions (MSCA Hybrid Intelligence to monitor, promote, and analyze transformations in good democracy practices).

### v3

@@ 7.4.6. Application of Large Language Models
With the emergence of Large Language Models (LLMs), the landscape of offensive language detection has rapidly shifted. Models such as GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), Fine-tuned LAnguage Net (FLAN) (Wei et al., 2021) have demonstrated impressive multilingual capabilities, and recent surveys have highlighted their growing importance (Zhao et al., 2023). Instruction-tuned LLMs with prompting can already perform competitive detection across monolingual and multilingual settings without the need for task-specific fine-tuning (Plaza-del arco et al., 2023). Prompt design strategies, such as incorporating task descriptions or input-label demonstrations, further enhance performance (Chen et al., 2022; Han and Tang, 2022; Roy et al., 2023). LLMs also show promise in addressing data scarcity by generating synthetic examples for low-resource languages (Hartvigsen et al., 2022), and in providing explanations that improve interpretability (Wang et al., 2023). Despite their advantages, LLMs are not a full replacement for prior methods. They often underperform specialised multilingual PLMs on fine-grained tasks when little prompting expertise is available, and they remain costly to deploy at scale. Traditional approaches, such as leveraging domain-specific lexicons (e.g., HurtLex) or designing language-family-specific PLMs (e.g., MuRIL, AraBERT), still work well for targeted low-resource or domain-specific scenarios, offering more efficiency and control.
Overall, while early work on LLMs in this area was scarce at the time of our survey design, the rapid pace of development since then underscores their growing importance. Future studies should explicitly and systematically compare LLM-based approaches with prior CLTL techniques to establish where LLMs offer clear advantages and where CLTL methods remain more effective. Another important research direction is to explore the complementarity between LLMs and CLTL approaches. For example, lexicon features, linguistic distance, and culturally-aware representations can be integrated as prompts or external knowledge sources to guide LLMs, while smaller multilingual PLMs can be lightweight alternatives in resource-constrained environments.
@@ 8. Conclusion
  • Cultural sensitivity: Offensive expressions are often context-dependent, and models trained on data from one cultural context may misclassify content in another, e.g., sarcastic expressions such as “well done” or idioms like calling someone a “clever fox” may be benign in one language but misinterpreted as neutral or offensive in another cultural context when translated. Culturally-aware annotation and evaluation with native speakers from target communities are necessary.
@@ Acknowledgements.
Aiqi Jiang is funded by China Scholarship Council (CSC Funding, No. 201908510140). Arkaitz Zubiaga acknowledges support from the European Union and UK Research and Innovation under Grant No. 101073351 as part of Marie Skłodowska-Curie Actions (MSCA Hybrid Intelligence to monitor, promote, and analyse transformations in good democracy practices).